Caspian Sea Drinks needs to raise $50.00 million by issuing additional shares of stock. If the market estimates CSD will pay a dividend of $2.66 next year, which will grow at 3.89% forever and the cost of equity to be 10.11%, then how many shares of stock must CSD sell?

Suppose the risk-free rate is 2.93% and an analyst assumes a market risk premium of 7.06%. Firm A just paid a dividend of $1.45 per share. The analyst estimates the β of Firm A to be 1.38 and estimates the dividend growth rate to be 4.26% forever. Firm A has 291.00 million shares outstanding. Firm B just paid a dividend of $1.96 per share. The analyst estimates the β of Firm B to be 0.75 and believes that dividends will grow at 2.48% forever. Firm B has 185.00 million shares outstanding. What is the value of Firm A?
Answer #1

1.Price of Share =Dividend Next year/(Cost of equity-growth) =2.66/(10.11%-3.89%) =42.7652733118971
Number of Shares to be sold to raise 50 million =50000000/42.7652733118971=1169172.93 or 1169173 shares


2.
cost of equity of A using CAPM model =Risk Free Rate+Beta*(Market Return-Risk Free Rate) =2.93%+1.38*7.06% =12.6728%
Price of Share A =Dividend Next year/(Cost of Equity-growth)=1.45*(1+4.26%)/(12.6728%-4.26%) =17.969879231647
Value of firm A =Number of Shares*Price of Share =291000000*17.969879231647=5,229,234,856.41